OpenStudy (anonymous):

solve each equation 5 raised to the 2x+1=25

OpenStudy (anonymous):

\[\LARGE 5^{2x+1}=25\] you can write \[\LARGE 25=5^2\] so...

OpenStudy (anonymous):

5^2 = 5^(2x + 1) 2 = 2x + 1 2x = 1 x = 1/2

OpenStudy (anonymous):

pratu's right :)